what is the solutions of this equation?
[4-x]+9=10

Answer by jim_thompson5910(35256)

Assuming those are absolute value brackets


|4-x|+9 = 10


|4-x| = 10-9


|4-x| = 1


4-x = 1 or 4-x = -1


-x = 1-4 or -x = -1-4


-x = -3 or -x = -5


x = 3 or x = 5